OpenStudy (anonymous):

-1/6(x-18)+1/2(x+2)=x+16

OpenStudy (anonymous):

-1/6x+3+1/2x+1=x+16 -1/6x+1/2x-x=16-4 -1/6x+3/6x-6/6x=12 -4/6x=12 6x=12/-4 6x=-3 x=-3/6 x=-1/2